year cash flow

0 -15000

1 +12000

2 +6000

3 -8000

4 +4000

5 +12000

use descartes and norstrom rules to define the characteristic of internal rate of return this investment

according to Descartes rules

according to Norstrom criteria

B) find external rate of return of the investment opportunity if the external investment %10
Answer #1
year cash flow Cumulative cash Flow
0 -15000 -15000
1 12000 -3000
2 6000 3000
3 -8000 -5000
4 4000 -1000
5 12000 11000
Number of sign changes 3 3
According toDescarte's Rule,number of internal rate of return values isalways less than or equal to number of sign changes
in the cash flow sequence
In this case there are three sign changes in cash flow sequence.Hence there can be three real positive values of InternalRate of Return
Norstrom's Rule is about sign changes in cumulative cash flow
According to Norstrom's Rule,there is only one internal rate of return value if cumulative cash flow that starts with negative and changes sign only once
In this case there are three sign changes in Cumulative cash flow sequence.HenceNorsstrom's rule is INCONLUSIVE
EXTERNAL   RATE OF RETURN
Extenal Invetment Rate=10%=0.1
Future Value of Cash Flow= (Cash flow)*(1.1^(5-N))
N= Year of Cash flow
N A B
year cash flow Future Value
1 12000 17569
2 6000 7986
3 -8000 -9680
4 4000 4400
5 12000 12000
SUM 32275
External Rate of Return=R
15000*((1+R)^5)=32275
(1+R)^5=32275/15000= 2.15168
1+R=(2.15168^(1/5)= 1.165616
R= 0.165616
External Rate Of Return = 0.165616
External Rate Of Return in percentage 16.56%